OpenStudy (anonymous):

400 cm3 of mercury at 0 degrees Celcius will expand to what volume at 50 degrees Celcius?

OpenStudy (anonymous):

oh. this one..

OpenStudy (anonymous):

what is the thermal expansion constant for mercury though?

OpenStudy (anonymous):

\[\Delta V=V_0(1+\gamma\Delta T)\] where delta V = change in volume in time delta T V0 = initial volume gamma = vometric thermal expantion constant for mercury

OpenStudy (anonymous):

@Brybro I am here..

OpenStudy (anonymous):

hey i actually figured this one out. can you help me with a different one please?

OpenStudy (anonymous):

go right ahead.

OpenStudy (anonymous):

20.00 cm of total space available. How long of a piece of brass at 20 degrees Celsius can be put there and still fit at 200 degrees Celsius?

OpenStudy (anonymous):

Brass has a linear expansion coefficient of 19X10^-6

OpenStudy (anonymous):

ok.. same formula. \(\Delta L=20\) if all units are: cm, and celcius, then \[\Delta L=L_0(1+\alpha \Delta T)\\ 20=L\times[1+19\times10^{-6}\times(200-20)] \]

OpenStudy (anonymous):

I can't understand that formula. lol

OpenStudy (anonymous):

oh.. ok. wait..